OpenStudy (anonymous):

prove sin theta tan theta = 1-cos squared theta/cos theta

OpenStudy (anonymous):

\[\sin \theta \tan \theta = \frac{1-\cos^2 \theta}{\cos \theta}\] \[LHS=\sin \theta \tan \theta =\sin \theta \times \frac{\sin \theta} {\cos \theta} =\frac{\sin^2 \theta} {\cos \theta}\] \[=\frac{1-\cos^2 \theta} {\cos \theta} =RHS\] Hence LHS=RHS i.e. \[\sin \theta \tan \theta = \frac{1-\cos^2 \theta}{\cos \theta}\] Thus proved. @shawnray

OpenStudy (anonymous):

Reasons: \[\tan \theta = \frac{\sin \theta}{\cos \theta} \] \[\sin^2 \theta = 1- \cos^2 \theta\]

OpenStudy (anonymous):

@shawnray

OpenStudy (anonymous):

thank you